***Please answer Question 1 AND Question 2 to receive positive raising. *** Please answer ALL.

Question 1
Question Help * P 8-31 (similar to) You need a particular piece of equipment for your production process. An equipment-leasing company has offered to lease the equipment to you for $10,500 per year if you sign a guaranteed 5-year lease (the lease is paid at the end of each year). The company would also maintain the equipment for you as part of the lease. Alternatively, you could buy and maintain the equipment yourself. The cash flows from doing so are listed below (the equipment has an economic life of 5 years). If your discount rate is 6.7%, what should you do? Year 1 -$2,200 Year 2 Year 3 $2,200 Year 4 Year 5 Year 0 2,200 2,200 2,200 0.700


Question 2

Question Help P8-36 (slmilar to) Fabulous Fabricators needs to decide how to allocate space in its production facility this year, It is considering the following contracts: NPV $2.01 million $0.95 million S1.46 million Use of Facility 100% 51% 49% Contract a. What are the profitability indexes of the projects? b. What should Fabulous Fabricators do?
Answer #1

Answer:

Question 1)

Given

Leasing cost per year L=$10500

interest rate r=6.7%

n=5 years

PV of leasing agreement =- L*(1-(1+r)^-n)/r=10500*(1-(1+6.7%)^-5)/6.7%=$43400.11 Eq 1

PV of if we buy equipment

PV =CF/(1+r)^t

where CF =cash flow

r=6.7%

t=Year in which cash flow occur

Year Cashflow PV of Cash flow
0 40700 40700.00
1 2200 2061.86
2 2200 1932.39
3 2200 1811.05
4 2200 1697.32
5 2200 1590.75
Total 49793.36

PV of if buy machine = Sum of PV of cash flow=$49793.36 Eq 2

From Equation 1 and 2 cost of leasing is we find that cost of leasing is lower than cost of buying so we will lease out the machine.